# Activation of entanglement in generalized entanglement swapping

### v2

@@ Appendix A Post-measurement state of (1,2) and (3,4) pairs
To get the post-measurement state of (1,2) and (3,4) for $n$-th outcome, we rewrite the Eq.5 in computational basis form. Every orthonormal basis $|\phi_{n\alpha}\rangle$ can be written in the following computational basis form $|\phi_{n\alpha}\rangle=\sum_{i,j=0}^{1}a_{n\alpha|ij}|ij\rangle$ where $\sum_{i,j=0}^{1}|a_{n\alpha|ij}|^{2}=1$ and $a_{n\alpha|ij}\in\mathbb{C}$.
@@
$=$ $\frac{1}{4p_{n}}\sum_{\alpha,\beta=1}^{4}\sum_{\begin{subarray}{c}i,j,k,l,\\ i^{\prime},j^{\prime},\\ k^{\prime},l^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}a^{*}_{n\alpha|ij}a_{n\alpha|kl}a_{n\beta|i^{\prime}j^{\prime}}a^{\ast}_{n\beta|k^{\prime}l^{\prime}}$
$\qquad\qquad\qquad\qquad|ij\rangle\langle i^{\prime}j^{\prime}|_{14}\otimes|kl\rangle\langle k^{\prime}l^{\prime}|_{23}$
@@
Tracing out the 3 and 4 qubits from Eq.32 gives $j=j^{\prime}$ and $l=l^{\prime}$. So, we get the post-measurement state of $n$-th outcome between (1,2) pair which is given by
@@
Similarly, tracing out 1 and 2 qubits from Eq.32 gives $i=i^{\prime}$ and $k=k^{\prime}$. The post-measurement state of (3,4) pair of $n$-th outcome is
@@
From the above Eq.A, and Eq.36 one can easily compute the post-measurement states of $n$-th outcome between different pairs (1,2), and (3,4) respectively after the $\mathcal{M}_{1}$ POVM measurement.

### v3

@@ Appendix A Post-measurement state of (1,2) and (3,4) pairs
To get the post-measurement state of (1,2) and (3,4) for $n$th outcome, we rewrite Eq.5 in computational basis form. Every orthonormal basis $|\phi_{n\alpha}\rangle$ can be written in the following computational basis form $|\phi_{n\alpha}\rangle=\sum_{i,j=0}^{1}a_{n\alpha|ij}|ij\rangle$ where $\sum_{i,j=0}^{1}|a_{n\alpha|ij}|^{2}=1$ and $a_{n\alpha|ij}\in\mathbb{C}$:
@@
$=$ $\frac{1}{4p_{n}}\sum_{\alpha,\beta=1}^{4}\sum_{\begin{subarray}{c}i,j,k,l,\\ i^{\prime},j^{\prime},\\ k^{\prime},l^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}a^{*}_{n\alpha|ij}a_{n\alpha|kl}a_{n\beta|i^{\prime}j^{\prime}}a^{\ast}_{n\beta|k^{\prime}l^{\prime}}|ij\rangle\langle i^{\prime}j^{\prime}|_{14}\otimes|kl\rangle\langle k^{\prime}l^{\prime}|_{23}$
@@
Tracing out the three and four qubits from Eq.32 gives $j=j^{\prime}$ and $l=l^{\prime}$. So, we get the post-measurement state of $n$-th outcome between (1,2) pair which is given by
@@
Similarly, tracing out one and two qubits from Eq.32 gives $i=i^{\prime}$ and $k=k^{\prime}$. The post-measurement state of (3,4) pair of $n$th outcome is
@@
From the Eq.A, and Eq.36 one can easily compute the post-measurement states of $n$-th outcome between different pairs (1,2), and (3,4) respectively after the $\mathcal{M}_{1}$ POVM measurement.